Question 7, chap 11, sect 10. part of 10 points A merry-go-round rotates at the rate of 0.11 rad/s with a(n) 85.3 kg man standing at a point 2.5 m from the axis of rotation. What is the new angular speed when the man walks to a point 0.71 m from the center? Assume that the merry-go-round is a solid 7.41 x 102 kg cylinder with a radius of 3.80 m. Answer in units of rad/s.
Answer #1

Here ,

let the final angular speed is wf

initial angular speed , wi = 0.11 rad/s

initial angular momentum = final angular momentum

(0.50 * 741 * 3.80^2 + 85.3 * 2.5^2) * 0.11 = (0.50 * 741 * 3.80^2 + 85.3 * 0.71^2) * wf

solving for wf

wf = 0.12 rad/s

the final angular speed is 0.12 rad/s


answered by: ANURANJAN SARSAM
Answer #2

Here ,

let the final angular speed is wf

initial angular speed , wi = 0.11 rad/s

initial angular momentum = final angular momentum

(0.50 * 741 * 3.80^2 + 85.3 * 2.5^2) * 0.11 = (0.50 * 741 * 3.80^2 + 85.3 * 0.71^2) * wf

solving for wf

wf = 0.12 rad/s

the final angular speed is 0.12 rad/s